# SwarnRaft: Leveraging Consensus for Robust Drone Swarm Coordination in GNSS-Degraded Environments

### v1

@@ Abstract
Unmanned aerial vehicle (UAV) swarms are increasingly used in critical applications such as aerial mapping, environmental monitoring, and autonomous delivery. However, the reliability of these systems is highly dependent on uninterrupted access to the Global Navigation Satellite Systems (GNSS) signals, which can be disrupted in real-world scenarios due to interference, environmental conditions, or adversarial attacks, causing disorientation, collision risks, and mission failure. This paper proposes SwarnRaft, a blockchain-inspired positioning and consensus framework for maintaining coordination and data integrity in UAV swarms operating under GNSS-denied conditions. SwarnRaft leverages the Raft consensus algorithm to enable distributed drones (nodes) to agree on state updates such as location and heading, even in the absence of GNSS signals for one or more nodes. In our prototype, each node uses GNSS and local sensing, and communicates over WiFi in a simulated swarm. Upon signal loss, consensus is used to reconstruct or verify the position of the failed node based on its last known state and trajectory. Our system demonstrates robustness in maintaining swarm coherence and fault tolerance through a lightweight, scalable communication model. This work offers a practical and secure foundation for decentralized drone operation in unpredictable environments.
@@ I Introduction
Rapid proliferation of autonomous swarms, including unmanned aerial vehicles (UAVs), ground robots, and maritime systems, has enabled transformative applications in domains such as logistics, infrastructure monitoring, agriculture, and disaster response [1, 2, 3]. These distributed systems rely heavily on precise coordination, predominantly facilitated by Global Navigation Satellite Systems (GNSS) such as GPS, GLONASS, Galileo, and BeiDou [4]. However, GNSS signals are often unreliable in real-world conditions, such as urban canyons, dense vegetation, or indoor and subterranean environments. To mitigate these limitations, systems frequently integrate Inertial Navigation Systems (INS) to improve robustness and continuity.
As swarm deployments scale in size and mission complexity, achieving reliable consensus among agents becomes a critical challenge, one that draws direct parallels to problems in distributed computing and blockchain technologies. In this context, swarms must maintain accurate and synchronized state information (e.g., position, velocity, heading) despite partial failures and environmental interference [5, 6].
@@
UAV swarms operate under unique constraints, including limited compute and energy resources, intermittent connectivity, and bounded-delay communication networks. Consequently, traditional blockchain consensus protocols such as Proof-of-Work and Proof-of-Stake are ill-suited due to their computational and energy overhead [14, 15]. Similarly, Byzantine Fault-Tolerant (BFT) protocols, e.g., PBFT, Tendermint, Exonum [16, 17, 18], assume adversarial behavior and incur unnecessary complexity for typical crash-fault scenarios in swarms. This creates a compelling opportunity to adapt crash-tolerant consensus protocols, such as Raft [19], which are designed for environments where nodes (the drones) may fail by crashing, but not by acting maliciously. Raft’s lightweight leader-based architecture is well suited to the resource-constrained and real-time nature of swarm robotics.
@@
This work bridges the fields of swarm robotics and distributed consensus, advancing resilience in GNSS-degraded environments. SwarmRaft builds on established consensus algorithms, demonstrating their applicability to UAV swarm coordination. Our approach maintains theoretical rigor while prioritizing practical deployability, showing significant improvements in both simulated and real-world scenarios.
@@ II Related Work
Decentralized coordination mechanisms for UAV swarms have gained increasing attention, particularly through consensus-based and blockchain-backed approaches. While leader–follower architectures have shown success in formation control [20], their reliance on central nodes introduces vulnerabilities to single points of failure. In contrast, distributed strategies offer greater resilience. Tariverdi et al. [21] present a formation control framework tailored to UAV dynamics, and Jia et al. [22] demonstrate decentralized estimation and coordination in quadcopter swarms.
In the context of Industrial IoT, Seo et al. [23] propose a co-design framework that jointly optimizes communication scheduling and consensus execution to meet strict latency requirements. Although their setting involves fixed infrastructure, their insights into delay-aware consensus coordination are applicable to mobile multi-agent systems. Similarly, Ilić et al. [24] introduce adaptive asynchronous gossip algorithms for heterogeneous sensor networks, offering resilience and scalability through weighted message exchange without requiring global synchronization. While their work focuses on distributed estimation in sensor-rich environments, the underlying principles of asynchronous fault-tolerant consensus motivate swarm-scale extensions.
Blockchain-based protocols add guarantees of trust and tamper resistance in decentralized coordination. P-Raft [25] optimizes consensus in consortium settings, and DTPBFT [26] dynamically adjusts trust levels in UAV networks. Yazdinejad et al. [27] propose a zone-based drone authentication framework using a Delegated Proof-of-Stake (DDPOS) mechanism to enable secure, low-latency inter-zone handovers in smart city deployments.
Comparative analyses of consensus algorithms such as Raft, PBFT, and PoW have focused on trade-offs in scalability, fault tolerance, and energy use [28, 29, 30, 31]. Surveys such as [32] highlight the challenges of adapting these mechanisms to real-time robotics, particularly under energy and processing constraints.
Research on fault-tolerant swarm behavior includes predictive failure mitigation [33], self-healing topologies [34], and GNSS spoofing resilience [35]. Robust navigation under sensor failures [36] and adaptive topology control [37] further emphasize the need for distributed, resilient solutions.
@@ III-A Overview of the Consensus Scheme
In modern UAV swarms, individual GNSS or INS readings alone cannot guarantee resilience against sensor faults or deliberate spoofing. Our consensus scheme, SwarmRaft, leverages peer-to-peer distance measurements, crash fault-tolerant communication consensus, and a Byzantine-resilient evaluation mechanism to detect and correct malicious or faulty position reports. At each time step $k$, the swarm determines the position of each UAV (node). To determine the positions, nodes communicate via Raft consensus and elect a leader for the step. The leader collects every node’s raw GNSS and INS fused position, tests if it is consistent with others, recovers it if required, and returns the results to the nodes.
This design achieves fault tolerance under the assumption that up to $f$ nodes’ sensors may be corrupted, so long as $n\geq 2f+1$, where $n$ is the total number of nodes in the swarm. Honest nodes never collude to mislead, and initial positions are trusted. The result is a distributed protocol that detects spoofing and significant sensor drift, locally filters out unreliable data, and recovers accurate positions through consensus based on neighbor estimates. Communication cost remains $O(n)$ for the leader and $O(1)$ for other nodes.
@@
    Sense: Each UAV measures GNSS and INS.
@@
    Inform: Raw fused position is sent to the leader (as a client’s transaction).
@@
    Recover: If a sensor is determined to be faulty, the leader replaces its data with the median of peer estimates.
@@ III-B System & Threat Model
Each UAV in our scheme is modeled as a fully-capable sensing and communication node. At each discrete time step $k$, node $i$ collects an absolute position measurement $\mathbf{z}^{\rm GNSS}_{i,k}$ from its onboard GNSS receiver; because GNSS alone can drift or be spoofed, the UAV also carries an INS that produces short-term motion increments $\Delta\mathbf{u}_{i,k}$ through accelerometers and gyroscopes. To bind these two modalities and detect inconsistent readings, every UAV is further equipped with a ranging sensor (for example, ultra-wideband or RSSI-based) that yields noisy inter-node distances $d_{ij,k}$ to each neighbor $j$. Finally, nodes share all measurements and voting messages over an authenticated broadcast channel protected by pre-distributed cryptographic keys, so that neither message forgery nor tampering can occur undetected.
@@ III-C Position Estimation
At each time step $k\in\mathbf{N}$, the UAV with index $i\in\{1,\dots,n\}$ obtains two complementary measurements of its state. First, the GNSS receiver produces a direct but noisy readout of the true position $\mathbf{x}_{i,k}$:
@@
where $R_{\rm GNSS}$ is the GNSS measurement covariance, and $\mathcal{N}(m,\sigma^{2})$ is a normal distribution with mathematical expectation $m$ and variance $\sigma^{2}$. Second, the INS dead-reckons from the previous time step by integrating acceleration and rotation increments $\Delta\mathbf{u}_{i,k}$, yielding
$\mathbf{x}^{\rm INS}_{i,k}=\mathbf{x}_{i,k-1}+\Delta\mathbf{u}_{i,k}+\mathbf{v}^{\rm INS}_{i,k},\quad\mathbf{v}^{\rm INS}_{i,k}\sim\mathcal{N}(0,R_{\rm INS}).$
@@
This geometric constraint links the true positions of nodes $i$ and $j$ and can reveal inconsistencies when one node’s GNSS or INS has been spoofed or drifted excessively.
III-D Fused Estimate at Peer $j$ for Node $i$
Peer $j$ combines the raw GNSS reading from $i$ with a range-based estimate to form a minimum-variance fusion. Denote the range-only estimate as
$\mathbf{x}^{\rm range,(j)}_{i,k}=\mathbf{x}^{\rm INS}_{j,k}+\frac{d_{ij,k}}{\|\mathbf{x}^{\rm INS}_{j,k}-\mathbf{x}_{i,k-1}\|}\bigl{(}\mathbf{x}_{i,k-1}-\mathbf{x}^{\rm INS}_{j,k}\bigr{)}.$
A linear fusion of the two estimates,
$\hat{\mathbf{x}}^{(j)}_{i,k}=\alpha\,\mathbf{z}^{\rm GNSS}_{i,k}+(1-\alpha)\,\mathbf{x}^{\rm range,(j)}_{i,k},$
achieves minimum mean-square error when the weight $\alpha$ equals the ratio of the range-estimate variance to the total variance. Specifically, if $\sigma^{2}_{\rm GNSS}$ and $\sigma^{2}_{\rm range}$ denote the variances of the GNSS and range-based estimates (obtainable from $R_{\rm GNSS}$ and linearization of the range constraint), then
$\alpha=\frac{\sigma^{2}_{\rm range}}{\sigma^{2}_{\rm GNSS}+\sigma^{2}_{\rm range}},\qquad 1-\alpha=\frac{\sigma^{2}_{\rm GNSS}}{\sigma^{2}_{\rm GNSS}+\sigma^{2}_{\rm range}}.$
With this choice, the fused estimate $\hat{\mathbf{x}}^{(j)}_{i,k}$ has variance equal to the harmonic sum of the individual variances, ensuring optimal use of both GNSS and inter-UAV ranging information.
@@ III-E Fault Detection
In this phase, the leader scrutinizes the consistency of a reported position by comparing the fused estimate against the original GNSS reading. A large discrepancy indicates that either the reporting node’s sensors or the peer’s range measurement may be faulty. To make this precise, we define a residual, apply a statistically-justified threshold, and then cast a binary vote.
The leader measures the consistency of node’s $i$ report with the data from node $j$ based on the received positions $\mathbf{z}_{i,k}$ and $\mathbf{z}_{j,k}$, it first computes its own fused estimate $\hat{\mathbf{x}}^{(j)}_{i,k}$ as described earlier. The residual
$e^{(j)}_{i,k}=\bigl{\|}\hat{\mathbf{x}}^{(j)}_{i,k}-\mathbf{z}_{i,k}\bigr{\|}.$
measures the Euclidean distance between the peer’s estimate and the raw GNSS reading. Under nominal (non-attacked) conditions, this residual is driven by zero-mean sensor noise from both GNSS and ranging, so it remains small with high probability.
To decide whether $e^{(j)}_{i,k}$ is abnormally large, we set a threshold
[…]
where $\mu_{e}$ and $\sigma_{e}$ are the mean and standard deviation of the residual under honest-sensor operation (obtainable by offline calibration). By the Gaussian tail bound, the probability that an honest residual exceeds $T$ is at most
[…]
Thus any $e^{(j)}_{i,k}>T$ is strong evidence of a sensor fault or spoofing.

### v2

# SwarmRaft: Leveraging Consensus for Robust Drone Swarm Coordination in GNSS-Degraded Environments

@@ Abstract
Unmanned aerial vehicle (UAV) swarms are increasingly used in critical applications such as aerial mapping, environmental monitoring, and autonomous delivery. However, the reliability of these systems is highly dependent on uninterrupted access to the Global Navigation Satellite Systems (GNSS) signals, which can be disrupted in real-world scenarios due to interference, environmental conditions, or adversarial attacks, causing disorientation, collision risks, and mission failure. This paper proposes SwarmRaft, a blockchain-inspired positioning and consensus framework for maintaining coordination and data integrity in UAV swarms operating under GNSS-denied conditions. SwarmRaft leverages the Raft consensus algorithm to enable distributed drones (nodes) to agree on state updates such as location and heading, even in the absence of GNSS signals for one or more nodes. In our prototype, each node uses GNSS and local sensing, and communicates over WiFi in a simulated swarm. Upon signal loss, consensus is used to reconstruct or verify the position of the failed node based on its last known state and trajectory. Our system demonstrates robustness in maintaining swarm coherence and fault tolerance through a lightweight, scalable communication model. This work offers a practical and secure foundation for decentralized drone operation in unpredictable environments.
@@ I Introduction
Rapid proliferation of autonomous swarms, including unmanned aerial vehicles (UAVs), ground robots, and maritime systems, has enabled transformative applications in domains such as logistics, infrastructure monitoring, agriculture, and disaster response [1, 2, 3]. These distributed systems rely heavily on precise coordination, predominantly facilitated by Global Navigation Satellite Systems (GNSS) such as GPS, GLONASS, Galileo, and BeiDou [4]. However, GNSS signals are often unreliable in real-world conditions, such as urban canyons, dense vegetation, or indoor and subterranean environments. Even with recent advances in noise-resilient GNSS acquisition techniques [5], the technology remains vulnerable to spoofing, jamming, and environmental interference. To mitigate these limitations, systems frequently integrate Inertial Navigation Systems (INS) to improve robustness and continuity. This inherent limitation of standalone GNSS or INS systems motivates the need for collaborative, consensus-driven approaches, such as SwarmRaft, that combine GNSS, INS, and peer-to-peer fusion to achieve robust and fault-tolerant state estimation in UAV swarms.
As swarm deployments scale in size and mission complexity, achieving reliable consensus among agents becomes a critical challenge, one that draws direct parallels to problems in distributed computing and blockchain technologies. In this context, swarms must maintain accurate and synchronized state information (e.g., position, velocity, heading) despite partial failures and environmental interference [6].
@@
UAV swarms operate under unique constraints, including limited compute and energy resources, intermittent connectivity, and bounded-delay communication networks. Consequently, traditional blockchain consensus protocols such as Proof-of-Work and Proof-of-Stake are ill-suited due to their computational and energy overhead [14, 15]. Similarly, Byzantine Fault-Tolerant (BFT) protocols, e.g., Practical Byzantine Fault Tolerance (PBFT), Tendermint, Exonum [16, 17, 18], assume adversarial behavior and incur unnecessary complexity for typical crash-fault scenarios in swarms. This creates a compelling opportunity to adapt crash-tolerant consensus protocols, such as Raft [19], which are designed for environments where nodes (the drones) may fail by crashing, but not by acting maliciously. Raft’s lightweight leader-based architecture is well suited to the resource-constrained and real-time nature of swarm robotics.
@@
Motivated by the real-world challenges UAV swarms face, such as GNSS spoofing, jamming, and sensor failure, we develop a lightweight, fault-tolerant coordination framework that maintains collective situational awareness even when individual agents are compromised. Unlike centralized or resource-intensive approaches, SwarmRaft leverages crash-tolerant consensus and sensor fusion to operate efficiently in contested or degraded environments. This work bridges swarm robotics and distributed consensus, demonstrating that established protocols like Raft can be adapted for real-time UAV coordination. Our design balances theoretical soundness with practical deployability, yielding substantial improvements in localization accuracy and swarm resilience across simulated and real-world scenarios. SwarmRaft demonstrates how lightweight consensus protocols can enhance resilience and coordination in UAV swarms operating under GNSS-degraded or adversarial conditions.
@@ II Related Work
Decentralized coordination mechanisms for UAV swarms have gained increasing attention, particularly through consensus-based and blockchain-backed approaches. While leader-follower architectures have shown success in formation control [20], their reliance on central nodes introduces vulnerabilities to single points of failure. In contrast, distributed strategies offer greater resilience. Tariverdi et al. [21] present a formation control framework tailored to UAV dynamics, and Jia et al. [22] demonstrate decentralized estimation and coordination in quadcopter swarms. Similarly, Zuo et al. [23] propose a voting-based scheme for leader election in lead–follow UAV swarms under constrained communication ranges, representing an earlier Raft-inspired approach tailored to leader selection rather than distributed fault-tolerant localization.
In the context of Industrial IoT (IIoT), Seo et al. [24] propose a co-design framework that jointly optimizes communication scheduling and consensus execution to meet strict latency requirements. Although their setting involves fixed infrastructure, their insights into delay-aware consensus coordination are applicable to mobile multi-agent systems. Similarly, Ilić et al. [25] introduce adaptive asynchronous gossip algorithms for heterogeneous sensor networks, offering resilience and scalability through weighted message exchange without requiring global synchronization. While their work focuses on distributed estimation in sensor-rich environments, the underlying principles of asynchronous fault-tolerant consensus motivate swarm-scale extensions.
Blockchain-based protocols add guarantees of trust and tamper resistance in decentralized coordination. P-Raft [26] optimizes consensus in consortium settings, and Dynamic Trust Practical Byzantine Fault Tolerance (DTPBFT) [27] dynamically adjusts trust levels in UAV networks. Yazdinejad et al. [28] propose a zone-based drone authentication framework using a Delegated Proof-of-Stake (DDPOS) mechanism to enable secure, low-latency inter-zone handovers in smart city deployments, and fairness-oriented multi-UAV communication frameworks for urgent tasks [29].
Comparative analyses of consensus algorithms, such as Raft, PBFT, and Proof-of-Work (PoW) have focused on trade-offs in scalability, fault tolerance, and energy use [30, 31, 32, 33] , and communication efficiency optimizations in UAV ad hoc networks [34]. Surveys such as [35] highlight the challenges of adapting these mechanisms to real-time robotics, particularly under energy and processing constraints.
Research on fault-tolerant swarm behavior includes predictive failure mitigation [36], self-healing topologies [37], and GNSS spoofing resilience [38], as well as UAV-aided localization via RF characteristics [39]. Robust navigation under sensor failures [40] and adaptive topology control [41] further emphasize the need for distributed, resilient solutions.
@@ III-A Overview of the Consensus Scheme
In modern UAV swarms, individual GNSS or INS readings alone cannot guarantee resilience against sensor faults or deliberate spoofing. Our consensus scheme, SwarmRaft, leverages peer-to-peer distance measurements, crash fault-tolerant communication consensus, and a Byzantine-resilient evaluation mechanism to detect and correct malicious or faulty position reports. At each time step $k$, the swarm determines the position of each UAV (node). To determine the positions, nodes communicate via Raft consensus and elect a leader for the step. The leader collects every node’s position estimate (derived from GNSS when available and maintained by INS otherwise), tests if it is consistent with others, recovers it if required, and returns the results to the nodes.
This design achieves fault tolerance under the assumption that up to $f$ nodes’ sensors may be corrupted, so long as $n\geq 2f+1$, where $n$ is the total number of nodes in the swarm. Honest nodes never collude to mislead, and initial positions are trusted. The result is a distributed protocol that detects spoofing and significant sensor drift, locally filters out unreliable data, and recovers accurate positions through consensus based on neighbor estimates.
@@
    Sense: Each UAV measures its position(GNSS/INS) and internode distances.
@@
    Inform: Measurement data is sent to the leader (as a client’s transaction).
@@
    Recover: If a sensor is determined to be faulty, the leader replaces its data with the position computed with peer estimates.
@@ III-B System & Threat Model
Each UAV in our scheme is modeled as a fully-capable sensing and communication node. At each discrete time step $k$, node $i$ collects an absolute position measurement $\mathbf{z}^{\rm GNSS}_{i,k}$ from its onboard GNSS receiver; because GNSS alone can drift or be spoofed, the UAV also carries an INS that produces short-term motion increments $\Delta\mathbf{u}_{i,k}$ through accelerometers and gyroscopes. To bind these two modalities and detect inconsistent readings, every UAV is further equipped with a ranging sensor, for example, ultra-wideband or Received Signal Strength Indicator (RSSI)-based) that yields noisy inter-node distances $d_{ij,k}$ to each neighbor $j$. Finally, nodes share all measurements and voting messages over an authenticated broadcast channel protected by pre-distributed cryptographic keys, so that neither message forgery nor tampering can occur undetected.
@@ III-C Position Estimation
TABLE I: Summary of Notation
Symbol Description
$x_{i,k}$ True position of drone $i$ at time $k$
$z_{i,k}^{GNSS}$ GNSS position measurement of drone $i$
$v_{i,k}^{GNSS}$ GNSS measurement noise
$x_{i,k}^{INS}$ Position estimated from INS
$v_{i,k}^{INS}$ INS sensor noise or drift
$d_{ij,k}$ Measured distance from drone $i$ to $j$
$\eta_{ij,k}$ Noise in range measurement $d_{ij,k}$
$R_{GNSS},R_{INS}$ Covariance matrices for GNSS and INS noise
$\sigma_{d}^{2}$ Variance of range noise
NN Number of drones in the swarm
$f$ Number of compromised/faulty drones
At each time step $k\in\mathbf{N}$, the UAV with index $i\in\{1,\dots,n\}$ obtains two complementary sources of information aboutmeasurements of its state. First, the GNSS receiver produces a direct but noisy readout of the true position $\mathbf{x}_{i,k}$:
@@
where $z_{i,k}^{\text{GNSS}}$ is the GNSS measurement, $v_{i,k}^{\text{GNSS}}\in\mathbb{R}^{3}$ is the GNSS measurement noise, and $R^{\text{GNSS}}\in\mathbb{R}^{3\times 3}$ is its covariance matrix. Here, $\mathcal{N}(\mu,\Sigma)$ denotes a multivariate Gaussian distribution with mean $\mu$ and covariance $\Sigma$.
Second, in the event of a GNSS outage, the UAV can fall back on its inertial navigation system. The INS predicts the next position $\mathbf{x}_{i,k}^{INS}\in\mathbb{R}^{3}$ based on the previous estimate $x_{i,k-1}$ and the inertial measurements $\Delta u_{i,k}$, which represent increments derived from accelerometer and gyroscope readings over the interval [k1,k][k-1,k]:
$x_{i,k}^{\text{INS}}=f\!\left(x_{i,k-1},\,\Delta u_{i,k}\right)+v_{i,k}^{\text{INS}},\quad v_{i,k}^{\text{INS}}\sim\mathcal{N}(0,R^{\text{INS}}),$
where $f(\cdot)$ denotes the INS state-propagation model, $v_{i,k}^{\text{INS}}\in\mathbb{R}^{3}$ is the accumulated INS error over one step, and $R^{\text{INS}}\in\mathbb{R}^{3\times 3}$ characterizes the associated covariance.
@@
where $d_{ij,k}$ is the measured distance from UAV $j$ to UAV $i$ at time step $k$, $x_{i,k}$ and $x_{j,k}$ are the true positions of UAVs $i$ and $j$, respectively, $\eta_{ij,k}$ is the measurement noise, and $\sigma_{d}^{2}$ is its variance.This geometric constraint links the true positions of nodes $i$ and $j$ and can reveal inconsistencies when one node’s GNSS add/or INS has been spoofed or drifted excessively.
III-D Fused Estimation through Probabilistic Localization
Each node $i$ measures its distances to all neighboring nodes and pairs this information with its own GNSS-derived position. Both the GNSS estimate and the set of computed inter-node distances are then transmitted to the leader node. Similarly, every other node in the network reports its GNSS position together with its locally computed distances to neighbors. Upon receiving the report from node $i$, the leader uses the GNSS location of $i$ and its distance measurements to approximate the possible locations of neighboring nodes. Since a single distance measurement corresponds to a circle (in 2D) or a sphere (in 3D) centered at node $i$, each neighbor’s location is constrained to lie within such a probabilistic region. By aggregating the distance-based regions provided by all nodes, the leader constructs an intersection of feasible regions that captures the most likely positions of every node. This process can be interpreted as a form of SwarmRaft’s fused estimation: node $i$’s GNSS information defines its own anchor position, while the distance measurements extend this knowledge to constrain neighboring nodes. The leader’s role is to integrate these partial constraints from all reporting nodes, producing a consistent, network-wide probabilistic localization of the fleet.
@@ III-E Fault Detection
In this phase, the leader evaluates the consistency of a node’s reported position by estimating its location from inter-node distances and the GNSS-based positions of its neighbors. Large discrepancies between a node’s self-reported GNSS position and the probabilistic regions derived from neighbors’ distance measurements indicate that either the reporting node’s sensors or the peers’ range measurements may be faulty.
To make this precise, we define a residual, apply a statistically justified threshold, and then cast a binary vote.
Specifically, for node $i$, the leader constructs feasible regions of its location from all non-faulty neighbor reports (as spheres/circles defined by measured distances). The reported GNSS position $\mathbf{z}_{i,k}$ is then compared against this feasible region. We define the residual as the minimum Euclidean distance between $\mathbf{z}_{i,k}$ and the feasible region boundary:
$e_{i,k}\;=\;\min_{\mathbf{x}\in\mathcal{R}_{i,k}}\bigl\|\mathbf{x}-\mathbf{z}_{i,k}\bigr\|,$
where $\mathcal{R}_{i,k}$ denotes the feasible region of node $i$’s position at time $k$ derived from verified neighbors.
Under nominal (non-attacked) conditions, this residual is primarily driven by sensor noise in GNSS and inter-drone ranging, so it remains small with high probability. To detect abnormal residuals, we define a threshold
$T\;=\;\mu_{e}\;+\;3\,\sigma_{e},$
where $\mu_{e}$ and $\sigma_{e}$ are obtained via offline calibration under honest sensor operation. By Gaussian tail bounds, the probability that an honest residual exceeds $T$ is less than 0.01. Any residual $e_{i,k}>T$ is thus treated as strong evidence of a fault or spoofing attempt.